# InfMasking: Unleashing Synergistic Information by Contrastive Multimodal Interactions

### v1

@@ 4.2.3 Experiments with 2 Modalities on Multimodal IMDb
Multimodal IMDb(MM-IMDb) arevalo2017gated is a real-world multimodal, multi-label dataset designed for movie genre classification. It poses two major challenges: significant class imbalance with genres such as comedy and drama dominating the label distribution, and substantial semantic discrepancy between visual (poster) and textual (plot’s description) modalities. Since genre prediction based on a single modality is often unreliable and the combination of both modalities can perform better arevalo2017gated , this underscores the need for effective modeling of multimodal interactions. We select both single-modal and multi-modal as baselines. For unimodal, we choose SimCLR (image-only) chen2020simple and CLIP (pretrained on image-text pairs) radford2021learning . For multimodal, we include CLIP, SLIP mu2022slip , and CoMM.
Tab. 4 summarizes the experimental results. Models trained on both modalities consistently outperformed their single-modality counterparts, further validating the importance of optimizing multimodal representation learning. InfMasking achieves the best overall performance, improving upon CoMM by 1.31% in weighted F1-score and 2.14% in macro F1-score. It is also worth noting that CLIP with its original public weights achieves 58.9% on weighted F1-score, outperforming CLIP fine-tuned on MM-IMDb (54.59%). This suggests that redundant information learning may not always benefit complex tasks such as genre prediction, which require complementary modality alignment dufumier2024align . These results demonstrate the robustness and generalizability of InfMasking in handling imbalanced, semantically heterogeneous, and multi-label multimodal classification tasks.
@@ 53. Answer: [N/A]
  54. Justification: We utilize the pre-processed datasets provided by MultiBench [30], which have been anonymized to safeguard personal privacy. For the MM-IMDb experiments, we employ publicly available open-source pre-trained models that adhere to established security guidelines.
@@ 60. Guidelines:
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
@@ Appendix A Experimental Details
Training details. We use AdamW [33] as the optimizer in all experiments. Detailed hyperparameters are listed in Tab. 6. Following [12] on MM-IMDb, we also use a cosine scheduler with final value $10^{-6}$ and a warmup over 10 epochs. And all models are trained for 100 epochs except for MM-IMDb which is trained for 70 epochs. All experiments are conducted on a single NVIDIA 4090 GPU with 24GB memory.
Table 6: Hyperparameters for InfMasking. Masking ratio is the ratio of masking for each masked view. The V&T CP and V&T EE are the contact prediction and end-effector position prediction tasks on Vision&Touch dataset respectively.
@@ B.2 Multibench
According to  [30], all datasets below have been pre-processed to ensure the removal of any personally identifiable information and to safeguard user privacy (some datasets don’t include any personal information, e.g. Vision&Touch and MM-IMDb).
@@
B.3 MM-IMDb
Multimodal IMDb (MM-IMDb) [2] is designed for movie genre prediction and comprises 25,959 films, each annotated with posters, plot summaries, genre labels, and metadata. Derived from the Movielens 20M dataset [16], this benchmark focuses on 23-way multi-label classification. In our experiments, we utilize the image modality (movie posters) and the text modality (plot summaries). While MM-IMDb is part of the Multibench benchmark [30], we present it separately in our experiments, as our model is trained directly on the raw data instead of relying on the pre-processed features offered by Multibench.
@@ Appendix C Broader Impact
Privacy and security. As discussed in Section 4.2.1, the datasets used in this study span multiple domains, including healthcare, sentiment analysis and multimedia. According to [30], all instances containing personal information have been rigorously anonymized and de-identified. And the Vision&Touch and MM-IMDb datasets do not contain any personally identifiable information. All experiments are conducted using irreversible, pre-extracted features, except for MM-IMDb, which is processed directly from raw data; no raw or reconstructable user data is accessed, thereby minimizing privacy risks.

### v2

@@ 4.2.3 Experiments with 2 Modalities on Multimodal IMDb
Multimodal IMDb($M$-IMDb) arevalo2017gated is a real-world multimodal, multi-label dataset designed for movie genre classification. It poses two major challenges: significant class imbalance with genres such as comedy and drama dominating the label distribution, and substantial semantic discrepancy between visual (poster) and textual (plot’s description) modalities. Since genre prediction based on a single modality is often unreliable and the combination of both modalities can perform better arevalo2017gated , this underscores the need for effective modeling of multimodal interactions. We select both single-modal and multi-modal as baselines. For unimodal, we choose SimCLR (image-only) chen2020simple and CLIP (pretrained on image-text pairs) radford2021learning . For multimodal, we include CLIP, SLIP mu2022slip , and CoMM.
Tab. 4 summarizes the experimental results. Models trained on both modalities consistently outperformed their single-modality counterparts, further validating the importance of optimizing multimodal representation learning. InfMasking achieves the best overall performance, improving upon CoMM by 1.31% in weighted F1-score and 2.14% in macro F1-score. It is also worth noting that CLIP with its original public weights achieves 58.9% on weighted F1-score, outperforming CLIP fine-tuned on $M$-IMDb (54.59%). This suggests that redundant information learning may not always benefit complex tasks such as genre prediction, which require complementary modality alignment dufumier2024align . These results demonstrate the robustness and generalizability of InfMasking in handling imbalanced, semantically heterogeneous, and multi-label multimodal classification tasks.
@@ 53. Answer: [N/A]
  54. Justification: We utilize the pre-processed datasets provided by MultiBench [30], which have been anonymized to safeguard personal privacy. For the $M$-IMDb experiments, we employ publicly available open-source pre-trained models that adhere to established security guidelines.
@@ 60. Guidelines:
    • The name of the license (e.g., $C$-BY 4.0) should be included for each asset.
@@ Appendix A Experimental Details
Training details. We use AdamW [33] as the optimizer in all experiments. Detailed hyperparameters are listed in Tab. 6. Following [12] on $M$-IMDb, we also use a cosine scheduler with final value $10^{-6}$ and a warmup over 10 epochs. And all models are trained for 100 epochs except for $M$-IMDb which is trained for 70 epochs. All experiments are conducted on a single NVIDIA 4090 GPU with 24GB memory.
Table 6: Hyperparameters for InfMasking. Masking ratio is the ratio of masking for each masked view. The V&T CP and V&T $E$ are the contact prediction and end-effector position prediction tasks on Vision&Touch dataset respectively.
@@ B.2 Multibench
According to  [30], all datasets below have been pre-processed to ensure the removal of any personally identifiable information and to safeguard user privacy (some datasets don’t include any personal information, e.g. Vision&Touch and $M$-IMDb).
@@
B.3 $M$-IMDb
Multimodal IMDb ($M$-IMDb) [2] is designed for movie genre prediction and comprises 25,959 films, each annotated with posters, plot summaries, genre labels, and metadata. Derived from the Movielens 20M dataset [16], this benchmark focuses on 23-way multi-label classification. In our experiments, we utilize the image modality (movie posters) and the text modality (plot summaries). While $M$-IMDb is part of the Multibench benchmark [30], we present it separately in our experiments, as our model is trained directly on the raw data instead of relying on the pre-processed features offered by Multibench.
@@ Appendix C Broader Impact
Privacy and security. As discussed in Section 4.2.1, the datasets used in this study span multiple domains, including healthcare, sentiment analysis and multimedia. According to [30], all instances containing personal information have been rigorously anonymized and de-identified. And the Vision&Touch and $M$-IMDb datasets do not contain any personally identifiable information. All experiments are conducted using irreversible, pre-extracted features, except for $M$-IMDb, which is processed directly from raw data; no raw or reconstructable user data is accessed, thereby minimizing privacy risks.